OpenStudy (lukecrayonz):

Solve for x e^(2x)+2e^(x)-8=0

OpenStudy (mimi_x3):

let u = e^x

OpenStudy (mimi_x3):

\[u^2 + 2u - 8=0\]

OpenStudy (mimi_x3):

i think that you're able to do it now

OpenStudy (anonymous):

Clever girl

OpenStudy (mimi_x3):

and dont forget to sub back in the \(u\) to solve for \(x\)

OpenStudy (anonymous):

\[ 2 e^x+e^{2 x}-8=\left(e^x-2\right) \left(e^x+4\right)=0 \]

OpenStudy (anonymous):

As @Mimi_x3 said, it should be easy from now on.

OpenStudy (mimi_x3):

yeah; your method is easier.

OpenStudy (lukecrayonz):

I got it, thanks:) How about: log(x-3)=2?

OpenStudy (mimi_x3):

i think that you should post it in another question..